OpenStudy (anonymous):

How do you find complex/imaginary roots of the cube root of -27 ?

OpenStudy (tkhunny):

Do you have DeMoivre?

OpenStudy (ikram002p):

\(\large x^3=-25\) \(\large x^3+25=0\) factor it

OpenStudy (ikram002p):

\(\large x^3=-27\) \(\large x^3+27=0\) factor it \(\large x^3+27=(x+3)(x^2-3x+9)\) \(\large (x+3)=0 , x=-3 \text{ (real root)}\) then usinf general formula for quiadratic equation solve \(\large x^2-3x+9=0\) \(\Huge \color{green}{x=\frac {-b \pm \sqrt{b^2-4ac}}{2a}}\) \(\Huge \color{green}{x=\frac {3 \pm \sqrt{9-36}}{2}}\) \(\Huge \color{green}{x=\frac {3 \pm \sqrt{-25}}{2}}\) \(\Huge \color{green}{x=\frac {3 \pm 5i}{2}}\) soo cubic imaginary root are \(\Huge \color{blue }{x=\frac {3 + 5i}{2}}\) and \(\Huge \color{blue }{x=\frac {3 - 5i}{2}}\)

OpenStudy (anonymous):

thanks, makes a lot of sense

OpenStudy (ikram002p):

ur wlc :)